OpenStudy (anonymous):

2x - y = 4 6x - 3y = 7 solve for x and y

OpenStudy (lgbasallote):

multiply the first equation by -3 -6x + 3y = -12 6x - 3y = 7 ========= 0 + 0 = -5 ========= 0 = 5 what do you think is the answer?

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

i dunno